### v2

@@ 3.1 Prompt Optimization
In prompt optimization Li et al. (2025b), we update the instruction $\phi$ of an LM (as opposed to a scaffolded LM). We prompt an LM optimizer with an objective $J(\phi)$ and (optionally) a textual feedback $f$.
In its simplest form, we provide query and response examples and prompt the LM to generate the corresponding instruction Khattab et al. (2023); Zhou et al. (2023).
@@
Provided with different initial instructions, other methods define the objective as generating variations. Each variation is evaluated on a validation set, and the top-performing versions are retained. In Zhou et al. (2023), the objective is to generate semantically similar instructions. Evolutionary methods Fernando et al. (2024); Guo et al. (2024) combine parts of multiple instructions and then introduce random changes; we call this process crossover and mutation. In OPRO Yang et al. (2024), we define the objective as generating instructions with a higher score using previous instructions, their scores, and a random sample of input-output pairs from the training data. To introduce fine-grained variations of instructions, SAMMO Schnabel & Neville (2024) uses a more granular representation of prompts by breaking them down into smaller sections, such as input formatting. This fine-grained structure allows SAMMO to adjust specific components, such as changing the input format from raw text to a structured format (e.g., XML). In a similar vain, ABO Ma et al. (2024) implements the instruction as step-by-step guidelines. Thus, the LM optimizer only updates the specific step that failed.
@@
Figure 7: Evolution of LM optimization approaches. LM optimization approaches have evolved from prompt optimization, then agents, and finally workflows. In all approaches, an LM acts as an optimizer for generating updates on $\phi$. Top: prompt optimization focuses on LMs as opposed to scaffolded LMs. The LM optimizer uses a single or multiple LM calls. Middle Experiential Learning extends to agents and focuses on extracting insight from tasks and refining tools. The LM optimizer uses a single LM call to update them from successful/unsuccessful execution traces. Bottom: AutoDiff Frameworks focus on workflows where computation is represented as a graph. Using textual feedback, the LM optimizer can optimize nodes of the graph (e.g., system prompts).
@@
Another line of work uses an LM to generate rich textual feedback $f$, and defines the objective as following the feedback $f$ Pryzant et al. (2023); Wang et al. (2024d); Chen et al. (2024c); Austin & Chartock (2024). Importantly, textual feedback is obtained from sample queries where the target LM fails. Additionally, these works also propose summarizing multiple pieces of feedback Chen et al. (2024c) or updates Austin & Chartock (2024).
@@
Self-discover Zhou et al. (2024a) focuses on the output structure of the instruction. The objective is to generate a JSON format with keys representing distinct reasoning steps. By filling in each key in an autoregressive manner during generation, it guides the target LM’s reasoning steps.
@@ 3.2 Experiential Learning
We borrow the term experiential learning from ExpeL Zhao et al. (2024a) to refer to the training of agents with language supervision. The LM optimizer considers an objective $J(\phi)$ (e.g., “improve set of tools according to feedback”, “extract reusable tool from execution trace”), user queries $\mathbf{q}$, execution trace $\tau$, optionally a textual feedback $f$, and generates the updated variable $\phi^{\prime}$. In particular, the execution trace $\tau$ of agents consists of a sequence of tool calls/responses and may end with a response to the user.
As discussed in Sec. 2.3.1, tools play an important role in agents. As shown in Fig. 8, $\phi$ is either the set of tools or insights that are provided to the agent’s system prompt. Specifically, prior works leverage successful and unsuccessful execution traces with different objectives.
@@ 3.2.1 Insights Extraction
ExpeL Zhao et al. (2024a) starts by collecting a dataset for training consisting of user queries $\mathbf{q}$ and corresponding execution trace $\tau$. Using a Reflexion agent Shinn et al. (2023), multiple execution traces may be sampled for a query $\mathbf{q}$. In particular, the Reflexion agent allows the agent to retry the task upon failing, which allows the collection of unsuccessful and successful trajectories $(\tau^{\text{succ}},\tau^{\text{fail}})$ on the same task instance $\mathbf{q}$. The LM optimizer refines a list of insights from successful and unsuccessful execution traces $(\tau^{\text{succ}},\tau^{\text{fail}})$ for a given query $\mathbf{q}$ or from $k$ successful trajectories $(\tau_{1}^{\text{succ}},...,\tau_{k}^{\text{succ}})$. At inference, those insights are fixed and provided to the scaffolded LM through its system prompt.
@@
Follow-up works on ExpeL Zhao et al. (2024a), such as AutoGuide Fu et al. (2024b), modify insights to be selectively added to the context at each step. Insights are specific to a state and contain a description of the state in which it is applicable. However, these methods require additional cost at inference since an LM summarizes the current state and then uses it to select the corresponding state-specific insight. AutoManual Chen et al. (2024b) modifies ExpeL Zhao et al. (2024a) to prevent distributional shift in offline RL, where at test-time the state distribution is different Wang et al. (2023a). However, their paper does not show strong evidence that training agents from language supervision suffers from distributional shift Wang et al. (2023a). Additionally, MSI Fu et al. (2024a) proposes to categorize each insight and retrieve the relevant one at test-time based on pre-defined rules.
@@
ExpeL Zhao et al. (2024a) also explores new settings such as transfer learning, where we use insights learned from a source dataset $\{\mathbf{q}^{\text{source}}\}$ and transfer it to a target dataset $\{\mathbf{q}^{\text{target}}\}$. In parametric training Pouyanfar et al. (2018), it consists of initializing the weights with the model trained on the source dataset. Then, finetuning these weights on the target dataset. In ExpeL, they initialize with insights extracted from a source dataset. Then, the LM optimizer’s objective is to generate refined insights to the target task provided with an execution trace (from a human). Note that the update is done in a single shot thanks to the LM’s ability to interpret intricate objectives. Additionally, they manually inspect execution traces from the agent with and without the insights. For instance, in ALFWorld Shridhar et al. (2020), one insight would be “when searching for an item, consider its nature and its typical usage.” This leads the agent to update its belief Andreas (2022) and avoid unnecessary actions to find objects. Furthermore, they find that another insight, “if an attempt to interact with an item fails […] consider alternative actions or locations,” leads to self-correction. These findings are important, they showcase that simple high-level insights influence the behavior of agents in some predictable way.
@@
Figure 8: Experiential learning focuses on tools and insights. Successful execution traces are used for extracting reusable tools Wang et al. (2025c) while unsuccessful execution traces are used to refine tools Zhang et al. (2024). Additionally, we can contrast successful and unsuccessful trajectories on the same query to extract insights Zhao et al. (2024a).
@@ 3.2.2 Tools Optimization
Tool optimization Zhang et al. (2024); Zhao et al. (2024c); Wang et al. (2025c) focuses on refining developer-defined code into more effective routines. The main challenge of tool optimization is ensuring that the skills are reusable and useful across episodes Wang et al. (2024a).
@@
A related term to tool, but slightly more general is skill Wang et al. (2024f). It refers to both a textual representation of a sequence of action and developer-defined code as tool. The term “skills” emphasizes two aspects of tool optimization. First, we want to make tools more efficient, thereby achieving tasks with fewer steps. Second, similarly to unsupervised RL (URL) Zhao et al. (2022); Eysenbach et al. (2019), we want to extract the set of behaviors possible in the environment. Drawing an analogy with principal component analysis, before we are given any tasks, we want to know what is the basis of behavior that exists in some environment. Similarly, these methods Liu et al. (2025b); Wang et al. (2024f) stores skills into a vector database and retrieve corresponding skill at inference. For instance, in CER Liu et al. (2025b) skills are represented as a sequence of tool calls with a high-level description of the skill. In AWM Wang et al. (2024f), they explore different textual representation of skills. We can also understand these methods as ‘curating’ in-context samples for retrieval of similar experiences in ExpeL Zhao et al. (2024a). However, the textual representation represents a challenge for formal verification Wang et al. (2025c).
@@
At each step, LearnAct Zhao et al. (2024c) and AgentOptimizer Zhang et al. (2024) sample an execution trace $\tau\sim\pi_{\phi}(\mathbf{q})$, and refine the set of tools for unsuccessful execution traces $\tau^{\text{fail}}$ LearnAct uses an execution feedback $f$ corresponding to execution errors. The LM optimizer’s objective is to address $f$ by either revising, adding a tool, or updating the system prompt with information about the tools. AgentOptimizer does not rely on the execution feedback $f$. The LM optimizer modifies the set of tools at each step by revising, removing, or adding a tool to solve the task that the model failed on. Moreover, they propose heuristics to account for the LM’s inability to correctly understand the failure modes and update tools. At each epoch, they evaluate the performance of the agent on the training data. If the performance is lower than the previous epoch, they roll back the set of tools. However, this is expensive and quickly leads to overfitting, where tools are not reusable.
@@
In contrast, ASI Wang et al. (2025c) extracts reusable tools from a successful execution trace. One important challenge in extracting a reusable tool is to make sure that it is both useful and bug-free. AgentOptimizer Zhang et al. (2024) checks the performance on the training set with and without the updated tools by simulating again on the same query. An LM is used to rewrite the execution trace with only the extracted tools. Then, the agent attempts to finish the execution. If it successfully solves the task, then the extract tool is kept. In contrast to transfer learning for insights, ASI explores domain adaptation Ovadia et al. (2019) in web browsing. They find that the agent can effectively reuse tools and adapt tools to the new domain.
@@
Similar to unsupervised RL Zhao et al. (2022), recent works propose to use an LM to both propose tasks and extract tools from rollouts on these tasks. In Voyager Wang et al. (2024a), the LM optimizer refines tools based on execution feedback and adds them to the skill set when successfully used to achieve a task. SkillWeaver Zheng et al. (2025) extracts tool directly from successful execution traces. Then, it produces test cases to check the long-term viability of the tool.
@@ 3.3 AutoDiff Frameworks
The term autodiff is inspired by deep learning frameworks Paszke et al. (2019) in the sense that it implements elementary operations for defining and optimizing workflows Yuksekgonul et al. (2025); Cheng et al. (2024); Zhou et al. (2024b); Yin (2024). The frameworks represent execution traces with a directed acyclic graph (DAG) where nodes correspond to variables or operations and edges represent dependencies. They roll out execution traces $\tau\sim\pi_{\phi}$ and obtain execution feedback $f$. Subsequently, prior works Yuksekgonul et al. (2025); Yin (2024); Zhou et al. (2024b); Wang et al. (2025a) use the chain rule to back-propagate $f$ through the graph. Each variable of the graph is updated by the LM optimizer according to its feedback. Alternatively, Trace Cheng et al. (2024) converts the DAG into text akin to a Python traceback. The LM optimizer updates variables in an autoregressive manner. We organize this section as follows. First, we cover graph abstraction to understand which variables of the workflow are optimized in different frameworks. Second, we focus on how execution traces are recorded and used for optimization.
@@ 3.3.1 Graph Abstraction
We refer to graph abstraction as the representation of nodes in the graph. This abstraction is foundational because it determines how workflows are formalized, traced, and subsequently optimized. There are two types of nodes in the graph: variables and operations. Variables correspond to non-parametric components of the workflow (e.g., system prompt) or intermediate values (result from operations). Operations correspond to computation done on variables, e.g., Python function, LM call. As an analogy with graph representation for neural networks, variables would be weights, and operations would be layers. However, different from neural networks, any node of the computational graph can be optimized, including operations Cheng et al. (2024). We distinguish between two lines of work, focusing on prompts or Python functions. It is important to note that in this survey, we focus on workflows within the scope of scaffolded LM. Thus, we do not consider applications of these frameworks for optimization instances (as opposed to workflows).
TextGrad Yuksekgonul et al. (2025) and other variants Yin (2024); Wang et al. (2025a); Zhou et al. (2024b) focus on textual variables. In particular, operations include concatenations and LM calls Yuksekgonul et al. (2025). Accordingly, trainable variables correspond to the system prompt or user input template of an LM call.
Trace Cheng et al. (2024) focuses on general-purpose workflows. Importantly, operations are not limited to LM calls, they can be arbitrary Python functions. In practice, we wrap Python functions with the Python decorator “bundle”. Additionally, they support the optimization of nodes corresponding to operations. In that sense, it goes beyond prompts to optimizing the workflow’s code.
@@
Figure 9: Graph execution and optimization using chain rule vs. formatting of execution trace. AutoDiff frameworks consist of three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update. Both approaches differ during their backward pass. Left: Using the chain rule, we compute a grad field on each variable corresponding to a feedback that addresses the feedback on its child variable. The LM optimizer updates each trainable variable following its grad field. Right: It performs a reverse topological sort on the computation graph starting from the output to format the prompt template of the optimizer. Together with the feedback, the LM optimizer generates trainable variables in an autoregressive manner.
@@ 3.3.2 Graph Execution and Optimization
User input Template $\phi$Input data $\mathbf{q}$concat($\phi$, $\mathbf{q}$)inputLM callresponseevaluate(output)feedback
@@
We refer to graph execution and optimization as the process of sampling an execution trace $\tau\sim\pi_{\phi}(\mathbf{q})$ and optimizing variables $\phi$ of the graph according to the execution feedback $f$. As shown in Fig. 9, it involves three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update. During forward computation, we collect an execution feedback $f$. Then, backward feedback propagation performs a reverse topological sort to either apply the chain rule on the execution feedback Yuksekgonul et al. (2025); Yin (2024); Wang et al. (2025a); Zhou et al. (2024b) or format a prompt template with the DAG Cheng et al. (2024). On the one hand, the chain rule enables structured credit assignment by propagating feedback through the graph in a fine-grained manner. On the other hand, encoding the entire execution trace as a single prompt template is more token-efficient.
@@
In the context of optimization in the language space, the chain rule does not refer exactly to the differentiation of composite functions in calculus. Instead, it points to the idea of when a transformation is applied to the output of another transformation. Then, the overall effect on an input can be understood by systematically combining the influence of each transformation in sequence. As shown in Fig. 10 and following the graph abstraction presented in Sec. 3.3.1. Operations, including LM call and concatenation, create nodes and attach a grad function. Most importantly, the grad function contains a prompt for an LM that informs about that operation’s role. Similarly, each variable has a role. These roles are provided in text to account for heterogeneous variables and operations in workflows. Accordingly, the backward feedback propagation calls the attached grad function at each operation in the reverse order of the DAG. Each gradient is a textual feedback with respect to the child node of that operation. For each trainable variable, the LM optimizer’s objective follows the feedback as well as the variable’s role. Follow-up work extends Textgrad Yuksekgonul et al. (2025) in the following ways. AdalFlow Yin (2024) extends operations to retrievers Lewis et al. (2020). GASO Wang et al. (2025a) modifies the gradient operation to take into account neighborhood nodes.

### v3

@@ 4.1 Prompt Optimization
In prompt optimization [110], we update the prompt of an LM (as opposed to a scaffolded LM). We prompt an LM optimizer with an objective (e.g., “improve this instruction according to the following examples of tasks where the model fails”) and (optionally) a textual feedback $f$. The updates made by the LM optimizer is evaluated on a validation set and the best one is selected. Thus, while these methods use a different objective in natural language, by heuristically selecting the best update on a validation set, they implicitly optimize for performance on the validation set. In its simplest form, we provide query and response examples and prompt the LM to generate the corresponding instruction [20, 21].
@@
Provided with different initial instructions, other methods define the objective as generating variations. Each variation is evaluated on a validation set, and the top-performing versions are retained. In [21], the objective is to generate semantically similar instructions. Evolutionary methods [111, 112] combine parts of multiple instructions and then introduce random changes; we call this process crossover and mutation. In OPRO [113], we define the objective as generating instructions with a higher score using previous instructions, their scores, and a random sample of input-output pairs from the training data. To introduce fine-grained variations of instructions, SAMMO [114] uses a more granular representation of prompts by breaking them down into smaller sections, such as input formatting. This fine-grained structure allows SAMMO to adjust specific components, such as changing the input format from raw text to a structured format (e.g., XML). In a similar vein, ABO [115] implements the instruction as step-by-step guidelines. Thus, the LM optimizer only updates the specific step that failed.
@@
Figure 8: Evolution of LM optimization approaches. LM optimization approaches have evolved from prompt optimization, then agents, and finally workflows. In all approaches, an LM acts as an optimizer for generating updates on prompts or tools. Top: prompt optimization focuses on LMs as opposed to scaffolded LMs. The LM optimizer uses a single or multiple LM calls. Middle: experiential learning extends to agents and focuses on extracting insight from tasks and storing them in the system message, or refining tools. The LM optimizer uses a single LM call to update them from successful or unsuccessful execution traces. Bottom: AutoDiff frameworks focus on workflows where computation is represented as a graph interleaving LM calls and developer-defined code as tools. Using textual feedback, the LM optimizer can optimize nodes of the graph (e.g., system prompts).
@@
Another line of work uses an LM to generate rich textual feedback $f$, and defines the objective as following the feedback $f$ [39, 38, 116, 117]. Importantly, textual feedback is obtained from sample queries where the target LM fails. Additionally, these works also propose summarizing multiple pieces of feedback [116] or updates [117].
@@
Self-discover [118] focuses on the output structure of the instruction. The objective is to generate a JSON format with keys representing distinct reasoning steps. By filling in each key in an autoregressive manner during generation, it guides the target LM’s reasoning steps.
@@ 4.2 Experiential Learning
We borrow the term experiential learning from ExpeL [10] to refer to the training of agents with language supervision. The LM optimizer considers an objective (e.g., “extract a general insight on why you failed this task”, “extract reusable tool from execution trace”), prompt $\mathbf{s}_{0}$, execution trace $\tau$, optionally a textual feedback $f$, and generates an insight or a tool. In particular, the execution trace $\tau$ of agents consists of a sequence of tool calls/responses and ends with a response to the user. As discussed in Sec. 3.3.1, tools play an important role in agents. As shown in Fig. 9, the objective corresponds to either improving the set of tools or extract insights that are provided to the agent’s system prompt. Specifically, prior works leverage successful and unsuccessful execution traces with different objectives.
@@ 4.2.1 Insights Extraction
ExpeL [10] starts by collecting a dataset for training consisting of prompt $\mathbf{s}_{0}$ and corresponding execution trace $\tau$. Using a Reflexion agent [108], multiple execution traces may be sampled for a prompt $\mathbf{s}_{0}$. In particular, the Reflexion agent allows the agent to retry the task upon failing, which allows the collection of unsuccessful and successful trajectories $(\tau^{\text{succ}},\tau^{\text{fail}})$ on the same prompt instance $\mathbf{s}_{0}$. The LM optimizer refines a list of insights from successful and unsuccessful execution traces $(\tau^{\text{succ}},\tau^{\text{fail}})$ for a given prompt $\mathbf{s}_{0}$ or from $k$ successful trajectories $(\tau_{1}^{\text{succ}},...,\tau_{k}^{\text{succ}})$. At inference, those insights are fixed and provided to the scaffolded LM through its system prompt.
@@
Follow-up works on ExpeL [10], such as AutoGuide [119], modify insights to be selectively added to the context at each step. Insights are specific to a state and contain a description of the state in which it is applicable. However, these methods require additional cost at inference since an LM summarizes the current state and then uses it to select the corresponding state-specific insight. AutoManual [120] modifies ExpeL [10] to prevent distributional shift in offline RL, where at test-time the state distribution is different [121]. However, their paper does not show strong evidence that training agents from language supervision suffers from distributional shift [121]. Additionally, MSI [122] proposes to categorize each insight and retrieve the relevant one at test-time based on pre-defined rules.
@@
ExpeL [10] also explores new settings such as transfer learning, where we use insights learned from a source dataset $\{\mathbf{s}_{0}^{\text{source}}\}$ and transfer it to a target dataset $\{\mathbf{s}_{0}^{\text{target}}\}$. In parametric training [123], it consists of initializing the weights with the model trained on the source dataset. Then, finetuning these weights on the target dataset. In ExpeL, they initialize with insights extracted from a source dataset. Then, the LM optimizer’s objective is to generate refined insights to the target task provided with an execution trace (from a human). Note that the update is done in a single shot thanks to the LM’s ability to interpret intricate objectives. Additionally, they manually inspect execution traces from the agent with and without the insights. For instance, in ALFWorld [124], one insight would be “when searching for an item, consider its nature and its typical usage.” This leads the agent to update its belief [125] and avoid unnecessary actions to find objects. Furthermore, they find that another insight, “if an attempt to interact with an item fails […] consider alternative actions or locations,” leads to self-correction. These findings are important, they showcase that simple high-level insights influence the behavior of agents in some predictable way.
@@
Figure 9: Experiential learning focuses on tools and insights. Successful execution traces are used for extracting reusable tools [11] while unsuccessful execution traces are used to refine tools [126]. Additionally, we can contrast successful and unsuccessful trajectories on the same task to extract insights [10].
@@ 4.2.2 Tools Optimization
Tool optimization [126, 127, 11] focuses on refining developer-defined code into more effective routines. The main challenge of tool optimization is ensuring that the skills are reusable and useful across episodes [128].
@@
A related term to tool, but slightly more general is skill [129]. It refers to both a textual representation of a sequence of action and developer-defined code as tool. The term “skills” emphasizes two aspects of tool optimization. First, we want to make tools more efficient, thereby achieving tasks with fewer steps. Second, similarly to unsupervised RL (URL) [130, 131], we want to extract the set of behaviors possible in the environment. Drawing an analogy with principal component analysis, before we are given any tasks, we want to know what is the basis of behavior that exists in some environment. Similarly, these methods [132, 129] stores skills into a vector database and retrieve corresponding skill at inference. For instance, in CER [132] skills are represented as a sequence of tool calls with a high-level description of the skill. In AWM [129], they explore different textual representation of skills. We can also understand these methods as ‘curating’ in-context samples for retrieval of similar experiences in ExpeL [10]. However, the textual representation represents a challenge for formal verification [11].
@@
At each step, LearnAct [127] and AgentOptimizer [126] sample an execution trace $\tau\sim\pi(\mathbf{s}_{0})$, and refine the set of tools for unsuccessful execution traces $\tau^{\text{fail}}$. LearnAct uses execution feedback $f$ corresponding to execution errors. The LM optimizer’s objective is to address $f$ by either revising, adding a tool, or updating the system prompt with information about the tools. AgentOptimizer does not rely on the execution feedback $f$. The LM optimizer modifies the set of tools at each step by revising, removing, or adding a tool to solve the task that the model failed on. Moreover, they propose heuristics to account for the LM’s inability to correctly understand the failure modes and update tools. At each epoch, they evaluate the performance of the agent on the training data. If the performance is lower than the previous epoch, they roll back the set of tools. However, this is expensive and quickly leads to overfitting, where tools are not reusable.
@@
In contrast, ASI [11] extracts reusable tools from a successful execution trace. One important challenge in extracting a reusable tool is to make sure that it is both useful and bug-free. AgentOptimizer [126] checks the performance on the training set with and without the updated tools by simulating again on the same query. An LM is used to rewrite the execution trace with only the extracted tools. Then, the agent attempts to finish the execution. If it successfully solves the task, then the extracted tool is kept. In contrast to transfer learning for insights, ASI explores domain adaptation [133] in web browsing. They find that the agent can effectively reuse tools and adapt tools to the new domain.
@@
Similar to unsupervised RL [130], recent works propose to use an LM to both propose tasks and extract tools from rollouts on these tasks. In Voyager [128], the LM optimizer refines tools based on execution feedback and adds them to the skill set when successfully used to achieve a task. SkillWeaver [134] extracts tools directly from successful execution traces. Then, it produces test cases to check the long-term viability of the tool.
@@ 4.3 AutoDiff Frameworks
The term autodiff is inspired by deep learning frameworks [135] in the sense that it implements elementary operations for defining and optimizing workflows [9, 27, 136, 28]. The frameworks represent execution traces with a directed acyclic graph (DAG) where nodes correspond to variables or operations and edges represent dependencies. They roll out execution traces $\tau\sim\pi$ and obtain execution feedback $f$. Subsequently, prior works [9, 28, 136, 137] use the chain rule to back-propagate $f$ through the graph. Each variable of the graph is updated by the LM optimizer according to its feedback. Alternatively, Trace [27] converts the DAG into text akin to a Python traceback. The LM optimizer updates variables in an autoregressive manner. We organize this section as follows. First, we cover graph abstraction to understand which variables of the workflow are optimized in different frameworks. Second, we focus on how execution traces are recorded and used for optimization.
@@ 4.3.1 Graph Abstraction
We refer to graph abstraction as the representation of nodes in the graph. This abstraction is foundational because it determines how workflows are formalized, traced, and subsequently optimized. There are two types of nodes in the graph: variables and operations. Variables correspond to non-parametric components of the workflow (e.g., system prompt) or intermediate values (result from operations). Operations correspond to computation done on variables, e.g., Python functions, LM calls. As an analogy with graph representation for neural networks, variables would be weights, and operations would be layers. However, different from neural networks, any node of the computational graph can be optimized, including operations [27]. We distinguish between two lines of work, focusing on prompts or Python functions. It is important to note that in this survey, we focus on workflows within the scope of scaffolded LM. Thus, we do not consider applications of these frameworks for optimization instances (as opposed to workflows).
TextGrad [9] and other variants [28, 137, 136] focus on textual variables. In particular, operations include concatenations and LM calls [9]. Accordingly, trainable variables correspond to the system prompt or user input template of an LM call.
Trace [27] focuses on general-purpose workflows. Importantly, operations are not limited to LM calls, they can be arbitrary Python functions. In practice, we wrap Python functions with the Python decorator “bundle”. Additionally, they support the optimization of nodes corresponding to operations. In that sense, it goes beyond prompts to optimizing the workflow’s developer-defined tool.
@@
Figure 10: Graph execution and optimization using chain rule vs. formatting of execution trace. AutoDiff frameworks consist of three key steps: (1) forward computation, (2) backward feedback propagation, (3) variable updates. Both approaches differ during their backward pass. Left: using the chain rule, we compute a grad field on each variable corresponding to a feedback that addresses the feedback on its child variable. The LM optimizer updates each trainable variable following its grad field. Right: it performs a reverse topological sort on the computation graph starting from the output to format the prompt template of the optimizer. Together with the feedback, the LM optimizer generates trainable variables in an autoregressive manner.
@@ 4.3.2 Graph Execution and Optimization
system messageuser messageconcat(system message, user message)inputLM callresponseevaluate(output)feedback
@@
We refer to graph execution and optimization as the process of sampling an execution trace $\tau\sim\pi(\mathbf{s}_{0})$ and optimizing variables of the graph according to the execution feedback $f$. As shown in Fig. 10, it involves three key steps: (1) forward computation, (2) backward feedback propagation, (3) variables update. During forward computation, we collect execution feedback $f$. Then, backward feedback propagation performs a reverse topological sort to either apply the chain rule on the execution feedback [9, 28, 137, 136] or format a prompt template with the DAG [27]. On the one hand, the chain rule enables structured credit assignment by propagating feedback through the graph in a fine-grained manner. On the other hand, encoding the entire execution trace as a single prompt template is more token-efficient.
@@
In the context of optimization in the language space, the chain rule does not refer exactly to the differentiation of composite functions in calculus. Instead, it points to the idea of when a transformation is applied to the output of another transformation. Then, the overall effect on an input can be understood by systematically combining the influence of each transformation in sequence. As shown in Fig. 11, and following the graph abstraction presented in Sec. 4.3.1, operations, including LM calls and concatenation, create nodes and attach a grad function. Most importantly, the grad function contains a prompt for an LM that informs about that operation’s role. Similarly, each variable has a role. These roles are provided in text to account for heterogeneous variables and operations in workflows. Accordingly, the backward feedback propagation calls the attached grad function at each operation in the reverse order of the DAG. Each gradient is a textual feedback with respect to the child node of that operation. For each trainable variable, the LM optimizer’s objective follows the feedback as well as the variable’s role. Follow-up work extends Textgrad [9] in the following ways. AdalFlow [28] extends operations to retrievers [77]. GASO [137] modifies the gradient operation to take into account neighboring nodes.